### v2

@@ 3.2. $m$-great Lie algebras and mixing
Whenever an $m$-tuple acts Diophantinely on every level, Proposition 3.5 applies. By (3.3), this is enough to conclude rapid mixing for the random walk they generate. Proposition 3.4 implies that the set of such tuples has full measure with respect to $\mu^{m}$. ∎
@@ 4. Nice Words
In this section, we prove Proposition 3.4. Our approach is to construct words in the semigroup $\langle F\rangle$ that differ by a nonzero element of $G^{(p)}$ for each $p$. In particular, we will find words that differ by a term that can be expressed as a $p$-fold bracket of elements of $\langle F\rangle$. Then, we will observe that a formula for this difference may be given as a polynomial in the exponential coordinates of the $g_{i}$. As long as the polynomials arising in this manner are linearly independent, we can apply Theorem 2.5 to conclude that, for a full measure set of walks, the words we construct act Diophantinely on the appropriate level, and thus we have nice words. This need for linear independence was the direct motivation for Definition 3.2. Once we have linear independence, Theorem 2.5 will guarantee the existence of a full measure of the set of $m$-tuples that act Diophanitnely and thus rapidly mix.
To begin, given $p\geq 1$, let $K_{p}\subset({\mathbb{Z}}^{m})^{(p+1)}$ be the set of all indices $\bf{k_{0},\ldots,k_{p}}$ such that $\ldots\bf{[k_{0}V,k_{1}V],\ldots],k_{p}V]}$ can be written of the form $\log(W_{1}W_{2}^{-1})\text{ mod }\mathfrak{g}^{(p+1)}$ for two words $W_{1}$ and $W_{2}$ of equal length in $\langle F\rangle$. This is the set of $p$-nested commutators that arise as the logarithm of the difference between two elements of our walk at some time. Thus $K_{p}$ can heuristically be seen as the set of elements of $\mathfrak{g}_{p}$ available to us in trying to get a self-cancellation between a character on level $p$ pushed forward by two words in our random walk. In constructing elements of $K_{p}$, the element $k_{ij}$ will count how many times $g_{j}$ appears in the $i^{th}$ step of the construction of the words differing by $\mathcal{H}_{p}(k,\alpha)$. For instance, in the setting of two generators $g_{1}$ and $g_{2}$, since
@@
for ${\bf k}_{0}=(1,0)$ and ${\bf k}_{1}=(0,1)$, we have that $((1,0),(0,1))\in K_{2}$. In terms of this set, our goal is to find an intersection between $K_{p}$ and the complement of $\mathcal{D}(\mathcal{H}_{m,p})$ i.e. a difference between two words in our walk with linearly independent polynomials. The following lemmas will demonstrate a method for finding elements of $K_{p}$ that will be rich enough that as long as $\mathcal{D}(\mathcal{H}_{m,p})\neq{\mathbb{R}}^{n_{p}}$, $K_{p}$ will not be contained in $\mathcal{D}(\mathcal{H}_{m,p})$ as we desire.
Lemma 4.1.
For any given $w_{-1},\ldots w_{p}\in\langle\mathcal{F}\rangle$ with $|w_{-1}|=|w_{0}|$, there exist words of equal length $W_{1}$ and $W_{2}$ such that
$W_{2}W_{1}^{-1}\in\mathfrak{g}_{p}.$
Proof.
Our idea is to inductively build words $a_{i}$ and $b_{i}$ in $\langle F\rangle$ such that $h_{i}:=a_{i}b_{i}^{-1}\in G^{(i)}$ and to do so with as much flexibility as possible. To begin with, we set
$a_{0}=w_{-1}\quad b_{0}=w_{0},$
and
@@
Since we took $|w_{-1}|=|w_{0}|$, this construction will always give us words $a_{i}$ and $b_{i}$ satisfying $|a_{i}|=|b_{i}|$. Observe that
@@
Since $a_{i-1}b_{i-1}^{-1}=h_{i-1}$, we can rewrite this as
$h_{i}=[h_{i-1},b_{i-1}w_{i}].$
Thus, $h_{1}$ can be expressed as a single bracket, and $h_{i}$ can be expressed as a bracket between $h_{i-1}$ and another group element, so inductively $h_{i}$ can always be expressed as an $i$-fold bracket. In particular, we get the formula
(4.1) $h_{p}=[\ldots[[a_{0}b_{0}^{-1},b_{0}w_{1}],b_{1}w_{2}],\ldots,b_{p-1}w_{p}],$
so $h_{p}\in\mathfrak{g}^{(p)}$ as desired. ∎
@@
For any $k$ satisfying $3k_{ij}<k_{i+1j}$, there exist words of equal length $W_{1},W_{2}$ such that
$\log(W_{2}W_{1}^{-1})=[\ldots\bf{[k_{0}V,k_{1}V],\ldots],k_{p}V]}\mod\mathfrak{g}^{(p+1)}.$
Proof.
By the previous Lemma, for any $w_{-1},\ldots w_{p}$ with $|w_{-1}|=|w_{0}|$, we can build words that differ by $h_{p}\in\mathfrak{g}^{(p)}$. Moreover, $h_{p}$ can be written as $a_{p}b_{p}^{-1}$ for two words of equal length in $\langle F\rangle$. If we can find ${\bf k_{0}},{\bf k_{1}},\ldots,{\bf k_{p}}$ such that
@@
then $({\bf k_{0}},{\bf k_{1}},\ldots,{\bf k_{p}})\in K_{p}$. Motivated by this, we take the logarithm of both sides of (4.1) to get
@@
Now, define a function ${\bf c}:\langle F,F^{-1}\rangle\rightarrow{\mathbb{Z}}^{m}$ such that ${\bf c}_{i}(W)$ is the number of times $g_{i}$ appears in the word $W$ with inverses counted as negatives. Define ${\bf k_{0}}={\bf c}(a_{0}b_{0}^{-1})$ and ${\bf k_{i}}={\bf c}(b_{i-1}w_{i})$. Applying BCH again, we have that $\log(b_{i-1}w_{i})=\sum_{i=1}^{m}{\bf k_{i}}V_{i}\mod\mathfrak{g}^{(1)}$. Applying this fact to (4.2) and using the fact that the $p$-fold bracket of a $\mathfrak{g}^{(1)}$ error is in $\mathfrak{g}^{p+1}$, we get
@@
Now, we have ${(\bf k_{0},\ldots,k_{p}})\in K_{p}$ which depends on our initial choice of $w_{-1},\ldots,w_{p}$. We wish to understand which sequences $\bf{k_{0},\ldots,k_{p}}$ are possible to get as a result of that choice. Note that $\bf{k_{0}}$ and $\bf{k_{1}}$ can be taken to be any vector of nonnegative integers simply by making an appropriate choice of $w_{-1},w_{0},$ and $w_{1}$. Meanwhile, for $i\geq 2$, we have ${\bf k_{i}}={\bf c}(b_{i-1})+{\bf c}(w_{i})$. Since by our choice of $w_{i}$ we can make ${\bf c}(w_{i})$ any nonnegative integer vector we like, this means we can choose any $\bf k_{i}$, subject to ${\bf k_{i}}\geq{\bf c}(b_{i-1})$. Since ${\bf c}(b_{i-1})=2{\bf c}(b_{i-2})+{\bf c}(w_{i-1})\leq 3{\bf k}_{i-1}$, if ${\bf k_{i}}\geq{\bf k_{i-1}}$, then ${(\bf k_{0},\ldots,k_{p}})\in K_{p}$. ∎
@@
By the previous lemma, $\{k\in({\mathbb{Z}}^{m})^{p+1}:3k_{ij}<k_{i+1j}\}$ is contained in $K_{p}$. It is clear that this set, when interpreted as a subset of ${\mathbb{R}}^{mp}$, is Zariski dense. It follows that $K_{p}$ is a Zariski dense set. ∎
@@
Next, by Theorem 2.5, since the coordinate polynomials of $\pi_{p}(W_{1}W_{2}^{-1})$ are linearly independent, for almost every value of the $\alpha_{ij}$, $\pi_{p}(W_{1}W_{2}^{-1})\in DC(\gamma,\tau).$ That means that for almost every $m$-tuple of $g_{i}$’s $\pi_{p}(W_{1}W_{2}^{-1})\in DC(\gamma,\tau).$ This implies the Proposition. ∎
@@ 5. Great Groups
Now that we have established that $m$-greatness implies rapid mixing of almost all random walks on $m$ elements, we wish to show that this condition is not overly restrictive. To that effect, we show that many special Lie algebras are 2-great, and that every step-$s$ Lie algebra is $s$-great. This will prove Theorems A and B. Finally, we present an example of a group that is not 2-great.
We now briefly discuss how we will approach the task of proving that a Lie algebra is $m$-great. Given a Lie algebra $\mathfrak{g}$, we will pick some concrete $k_{p}$ for which we will show that $\mathcal{H}_{m,p}(k_{p},\alpha)$ has linearly independent polynomials i.e. $k_{p}\notin\mathcal{D}(\mathcal{H}_{m,p})$. To show this linear independence, we let $\{P_{1},\ldots,P_{n_{p}}\}$ be the set of $n_{p}$ coordinate polynomials associated to the $X_{i}^{(p)}$. If there is a monomial $M_{i}=\alpha_{i_{0}j_{0}}\ldots\alpha_{i_{p}j_{p}}$ with nonzero coefficient in $P_{i}$ but zero coefficient in each $P_{j}$ when $j\neq i$, we call $M_{i}$ a unique monomial of $P_{i}$. If each $P_{i}$ has a unique monomial, this is sufficient to conclude that the $P_{i}$ form a linearly independent set. By finding a unique monomial of each of the $P_{i}$, we can show that the $P_{i}$ form a linearly independent set. Thus, $\bar{k}_{p}\notin\mathcal{D}(\mathcal{H}_{2,p})$ and $\mathcal{H}_{2,p}$ is non-degenerate. From there, we will conclude 2-greatness.
@@
Here we prove that all quasi-abelian, triangular and step-3 or lower Lie algebras are 2-great. Throughout this section, we shall denote $\bar{k}_{p}=((1,0),(0,1)^{p})$.
@@
and let $P_{1},\ldots,P_{n_{p}}$ be the coordinate polynomials in $\alpha$ of $\mathcal{H}_{2,p}(\bar{k}_{p},\alpha)$. We will show that if $P_{i}$ is the polynomial corresponding to the $Y_{i,p}$ coordinate then $\alpha_{0,i}\beta_{0}^{p}$ is a unique monomial of $P_{i}$. This is because the coefficient of $\alpha_{0,i}\beta_{0}^{p}$ in $\ldots[V_{1},V_{2}],\ldots V_{2}]$ is exactly $\ldots[[Y_{0,i},X],X],\ldots,X]=Y_{i,p}$. This means that each $P_{i}$ has a unique monomial, so they form a linearly independent set of polynomials. Thus $\bar{k}\notin\mathcal{D}(\mathcal{H}_{2,p})$, so $\mathcal{H}_{2,p}$ is non-degenerate. Since $p$ is arbitrary, any quasi-abelian Lie algebra is 2-great. ∎
@@
For any $s\in{\mathbb{N}}$, ${\mathbf{t}}_{s}$ is 2-great.
@@
Let $P_{1},\ldots,P_{n_{p}}$ be the coordinate polynomials in $\alpha$ of $\mathcal{H}_{2,p}(\bar{k}_{p},\alpha)$. We will show that if $P_{i}$ is the polynomial corresponding to the $E_{i,i+p+1}$ coordinate, then $\alpha_{i}\beta_{i+1},\ldots\beta_{i+p}$ is a unique monomial of $P_{i}$. This is because the coefficient of $\alpha_{i}\beta_{i+1},\ldots\beta_{i+p}$ in $\ldots[V_{1},V_{2}],\ldots V_{2}]$ is exactly $\ldots[E_{i,i+1},E_{i+1,i+2}],\ldots,E_{i+p,i+p+1}]=E_{i,i+p+1}$. This means that each $P_{i}$ has a unique monomial, so they form a linearly independent set of polynomials. Thus, $\bar{k}\notin\mathcal{D}(\mathcal{H}_{2,p})$ and so $\mathcal{H}_{2,p}$ is non-degenerate. Since $p$ is arbitrary, any triangular Lie algebra is 2-great. ∎
@@
We compute $\mathcal{H}_{2,0}$ and $\mathcal{H}_{2,1}$ for arbitrary groups. If $V_{1}=\sum_{i}\alpha_{i}X_{i}$, $V_{2}=\sum_{i}\beta_{i}X_{i}$, we have that $\mathcal{H}_{2,0}((1,0),\alpha)=V_{1}$ which clearly has linearly independent polynomials for coordinates. Thus $\mathcal{H}_{2,0}$ is always non-degenerate. Now, if $\mathfrak{g}$ is non-abelian, let $X^{(2)}_{j}=[X_{i},X_{i^{\prime}}]$. In $\mathcal{H}_{2,1}$, the monomial $\alpha_{i}\beta_{j}$ only appears as a coefficient of $X^{(2)}_{j}$. Thus, for any non-abelian Lie algebra $\mathcal{H}_{2,1}$ is non-degenerate. From this we conclude the Lemma. ∎
Now, we show that step-3 Lie algebras are also necessarily 2-great. This is the best we can do in terms of step as there exist step-4 Lie algebras that are not 2-great as we shall see. Our proof technique will be similar to the proofs of $2$-greatness that we have already presented in that we begin by fixing some choice of $k$ which we show to be in the complement of $\mathcal{D}(\mathcal{H}_{2,2})$. It will differ in that we will show that the existence of a linear dependence between the polynomials of $\mathcal{H}_{2,2}$ contradicts the fact that $\mathfrak{g}$ is step 3 rather than directly showing linear independence by identifying unique monomials of each polynomial.
@@
Since the equation above holds for arbitrary $i_{1},i_{2},i_{3}\in I$, this implies that $\lambda\!\!=\!\!0$, a contradiction. Thus, $\mathcal{H}_{2,2}$ is non-degenerate. The fact that $\mathcal{H}_{2,0}$ and $\mathcal{H}_{2,1}$ are non-degenerate follows from Lemma 5.3, so any step-3 (or lower) Lie algebra is 2-great. ∎

### v3

@@ 3.2. $m$-great Lie algebras and mixing
Whenever an $m$-tuple acts Diophantinely on every level, Proposition 3.5 applies. By (3.3), this is implies rapid mixing for the random walk they generate. By Proposition 3.4 the set of such tuples has full measure with respect to $\mu^{m}$. ∎
@@ 4. Nice Words
In this section, we prove Proposition 3.4. Our approach is to construct words in the semigroup $\langle F\rangle$ that differ by a nonzero element of $G^{(p)}$ for each $p$. In particular, we will find words that differ by a term that can be expressed as a $p$-fold bracket of elements of $\langle F\rangle$. Then, we will observe that a formula for this difference may be given as a polynomial in the exponential coordinates of the $g_{i}$. As long as the polynomials arising in this manner are linearly independent Theorem 2.5 implies that, for a full measure set of walks, the words we construct act Diophantinely on the appropriate level, and thus are nice. This need for linear independence was the direct motivation for Definition 3.2. Once we have linear independence, Theorem 2.5 guarantees the existence of a full measure of the set of $m$-tuples that act Diophanitnely and thus rapidly mix.
To begin, given $p\geq 1$, let $K_{p}\subset({\mathbb{Z}}^{m})^{(p+1)}$ be the set of all indices $\bf{k_{0},\ldots,k_{p}}$ such that $\ldots\bf{[k_{0}V,k_{1}V],\ldots],k_{p}V]}$ can be written of the form $\log(W_{1}W_{2}^{-1})\text{ mod }\mathfrak{g}^{(p+1)}$ for two words $W_{1}$ and $W_{2}$ of equal length in $\langle F\rangle$. This is the set of $p$-nested commutators that arise as the logarithm of the difference between two elements of our walk at some time. Thus $K_{p}$ can heuristically be seen as the set of elements of $\mathfrak{g}_{p}$ available to us in trying to get a self-cancellation between an element of $H_{\lambda}$ pushed forward by two words in our random walk when $\lambda\in\Lambda_{p}-\{0\}$. In constructing elements of $K_{p}$, the element $k_{ij}$ will count how many times $g_{j}$ appears in the $i^{th}$ step of the construction of the words differing by $\mathcal{H}_{p}(k,\alpha)$. For instance, in the setting of two generators $g_{1}$ and $g_{2}$, since
@@
for ${\bf k}_{0}=(1,0)$ and ${\bf k}_{1}=(0,1)$, we have that $((1,0),(0,1))\in K_{2}$. In terms of this set, our goal is to find an intersection between $K_{p}$ and the complement of $\mathcal{D}(\mathcal{H}_{m,p})$ i.e. a difference between two words in our walk with linearly independent polynomials. The following lemmas provide a method for finding elements of $K_{p}$ that will be rich enough so that as long as $\mathcal{D}(\mathcal{H}_{m,p})\neq{\mathbb{R}}^{n_{p}}$, $K_{p}$ will not be contained in $\mathcal{D}(\mathcal{H}_{m,p})$ as we desire.
We now describe a method to build inductively words $a_{i}$ and $b_{i}$ in $\langle F\rangle$ such that $h_{i}:=a_{i}b_{i}^{-1}\in G^{(i)}$ and to do so with as much flexibility as possible. In particular, given any choice of $w_{-1},\ldots w_{p}\in\langle F\rangle$ with $|w_{-1}|=|w_{0}|$, we can build the sequences $a_{i}$ and $b_{i}$ in $\langle w_{-1},w_{0},\dots,w_{p}\rangle$ To begin with, we set $a_{0}=w_{-1},b_{0}=w_{0},$ and
@@
Notice that $b_{i}$ will always be a concatenation of the $w_{-1},\ldots,w_{i}$. Since $|w_{-1}|=|w_{0}|$, this construction will always give us words $a_{i}$ and $b_{i}$ satisfying $|a_{i}|=|b_{i}|$. Observe that
@@
Since $a_{i-1}b_{i-1}^{-1}=h_{i-1}$, we can rewrite this as $h_{i}=[h_{i-1},b_{i-1}w_{i}].$ Thus, $h_{1}$ can be expressed as a single bracket, and $h_{i}$ can be expressed as a bracket between $h_{i-1}$ and another group element, so inductively $h_{i}$ can always be expressed as an $i$-fold bracket. In particular, we get the formula
(4.1) $h_{p}=[\ldots[[a_{0}b_{0}^{-1},b_{0}w_{1}],b_{1}w_{2}],\ldots,b_{p-1}w_{p}].$
We now show that this procedure produces a large set of group elements.
@@
Any multi-index ${\mathbf{k}}=({\mathbf{k}}_{0},\ldots,{\mathbf{k}}_{p})\in({\mathbb{Z}}^{m})^{(p+1)}$, ${\mathbf{k}}_{i}=(k_{i1},\dots,k_{im})$ satisfying $k_{(i+1)j}>3k_{ij}$, is contained in $K_{p}$.
If we can find ${\bf k_{0}},{\bf k_{1}},\ldots,{\bf k_{p}}$ such that $h_{p}$ in (4.1) satisfies
@@
then $({\bf k_{0}},{\bf k_{1}},\ldots,{\bf k_{p}})\in K_{p}$. Taking the logarithm of both sides of (4.1) we get
@@
Define a function ${\bf c}:\langle F,F^{-1}\rangle\rightarrow{\mathbb{Z}}^{m}$ such that ${\bf c}_{i}(W)$ is the number of times $g_{i}$ appears in the word $W$ with inverses counted as negatives. Define ${\bf k_{0}}={\bf c}(a_{0}b_{0}^{-1})$ and ${\bf k_{i}}={\bf c}(b_{i-1}w_{i})$. Applying BCH again, we have that $\log(b_{i-1}w_{i})={\bf k_{i}V}\mod\mathfrak{g}^{(1)}$. Applying this fact to (4.2) and using the fact that the $p$-fold bracket of a $\mathfrak{g}^{(1)}$ error is in $\mathfrak{g}^{p+1}$, we get
@@
Now, we have ${(\bf k_{0},\ldots,k_{p}})\in K_{p}$ which depends on our initial choice of $w_{-1},\ldots,w_{p}$. We wish to understand which sequences $\bf{k_{0},\ldots,k_{p}}$ are possible to get as a result of that choice. Note that $\bf{k_{0}}$ and $\bf{k_{1}}$ can be taken to be any vector of nonnegative integers simply by making an appropriate choice of $w_{-1},w_{0},$ and $w_{1}$. Meanwhile, for $i\geq 2$, we have ${\bf k_{i}}={\bf c}(b_{i-1})+{\bf c}(w_{i})$. Since by our choice of $w_{i}$ we can make ${\bf c}(w_{i})$ any nonnegative integer vector we like, this means we can choose any $\bf k_{i}$, subject to ${\bf k_{i}}\geq{\bf c}(b_{i-1})$. Since ${\bf c}(b_{i-1})=2{\bf c}(b_{i-2})+{\bf c}(w_{i-1})\leq 3{\bf k}_{i-1}$, if ${\bf k_{i}}\geq 3{\bf k_{i-1}}$, then ${(\bf k_{0},\ldots,k_{p}})\in K_{p}$.
@@
By the previous lemma, $\{{\textbf{k}}\in({\mathbb{Z}}^{m})^{p+1}:k_{(i+1)j}>3k_{ij}\}$ is contained in $K_{p}$. It is clear that this set, is Zariski dense in ${\mathbb{R}}^{m(p+1)}.$ It follows that $K_{p}$ is a Zariski dense set. ∎
@@
Next, by Theorem 2.5, since the coordinate polynomials of $\pi_{p}(W_{1}W_{2}^{-1})$ are linearly independent, for almost every value of the $\alpha_{ij}$, $\pi_{p}(W_{1}W_{2}^{-1})\in DC$. That means that for almost every $m$-tuple of $g_{i}$’s $\pi_{p}(W_{1}W_{2}^{-1})\in DC$. This implies the proposition. ∎
@@ 5. Great Groups
Now that we have established that $m$-greatness implies rapid mixing of almost all random walks on $m$ elements, we show that this condition is not overly restrictive. To that effect, we establish that many special Lie algebras are 2-great, and that every step-$s$ Lie algebra is $s$-great. This will prove Theorems A and B. Finally, we present an example of a group that is not 2-great.
We now briefly discuss our approach for proving that a Lie algebra is $m$-great. Given a Lie algebra $\mathfrak{g}$, we will pick some concrete $\bar{k}_{p}$ for which we will show that $\mathcal{H}_{m,p}(\bar{k}_{p},\alpha)$ has linearly independent polynomials i.e. $\bar{k}_{p}\notin\mathcal{D}(\mathcal{H}_{m,p})$. To show this linear independence, we let $\{P_{1},\ldots,P_{n_{p}}\}$ be the set of $n_{p}$ coordinate polynomials associated to the $X_{i}^{(p)}$. If there is a monomial $M_{i}=\alpha_{i_{0}j_{0}}\ldots\alpha_{i_{p}j_{p}}$ with nonzero coefficient in $P_{i}$ but zero coefficient in each $P_{j}$ when $j\neq i$, we call $M_{i}$ a unique monomial of $P_{i}$. By finding a unique monomial of each of the $P_{i}$, we can show that the $P_{i}$ form a linearly independent set. Thus, $\bar{k}_{p}\notin\mathcal{D}(\mathcal{H}_{2,p})$ and $\mathcal{H}_{2,p}$ is non-degenerate. From there, we will conclude 2-greatness.
@@
Here we prove that all quasi-abelian, triangular and step-3 or lower Lie algebras are 2-great. Throughout this section, we denote $\bar{k}_{p}=((1,0),(0,1)^{p})$.
@@
and let $P_{1},\ldots,P_{n_{p}}$ be the coordinate polynomials in $\alpha,\beta$ of $\mathcal{H}_{2,p}$. We will show that if $P_{i}$ is the polynomial corresponding to the $Y_{i,p}$ coordinate then $\alpha_{0,i}\beta_{0}^{p}$ is a unique monomial of $P_{i}$. This is because the coefficient of $\alpha_{0,i}\beta_{0}^{p}$ in $\ldots[V_{1},V_{2}],\ldots V_{2}]$ is exactly $\ldots[[Y_{0,i},X],X],\ldots,X]=Y_{i,p}$. This means that each $P_{i}$ has a unique monomial, so they form a linearly independent set of polynomials. Thus $\bar{k}\notin\mathcal{D}(\mathcal{H}_{2,p})$, so $\mathcal{H}_{2,p}$ is non-degenerate. Since $p$ is arbitrary, any quasi-abelian Lie algebra is 2-great. ∎
@@
For any $s\in{\mathbb{N}}$, the triangular algebra ${\mathbf{t}}_{s}$ is 2-great.
@@
Let $P_{1},\ldots,P_{n_{p}}$ be the coordinate polynomials in $\alpha,\beta$ of $\mathcal{H}_{2,p}$. We will show that if $P_{i}$ is the polynomial corresponding to the $E_{i,i+p+1}$ coordinate, then $\alpha_{i}\beta_{i+1},\ldots\beta_{i+p}$ is a unique monomial of $P_{i}$. Indeed the coefficient of $\alpha_{i}\beta_{i+1},\ldots\beta_{i+p}$ in $\ldots[V_{1},V_{2}],\ldots V_{2}]$ is exactly $\ldots[E_{i,i+1},E_{i+1,i+2}],\ldots,E_{i+p,i+p+1}]=E_{i,i+p+1}$. This means that each $P_{i}$ has a unique monomial, so they form a linearly independent set of polynomials. Thus, $\bar{k}\notin\mathcal{D}(\mathcal{H}_{2,p})$ and so $\mathcal{H}_{2,p}$ is non-degenerate. Since $p$ is arbitrary, any triangular Lie algebra is 2-great. ∎
@@
We compute $\mathcal{H}_{2,0}$ and $\mathcal{H}_{2,1}$ for arbitrary groups. If $V_{1}=\sum_{i}\alpha_{i}X_{i}$, $V_{2}=\sum_{i}\beta_{i}X_{i}$, we have that $\mathcal{H}_{2,0}((1,0),\alpha)=V_{1}$ which clearly has linearly independent polynomials for coordinates. Thus $\mathcal{H}_{2,0}$ is always non-degenerate. Now, if $\mathfrak{g}$ is non-abelian, let $X^{(1)}\in\mathfrak{g}^{(1)}$ be a nonzero vector of the form $X^{(1)}=[X_{i},X_{j}]$ for some $i$ and $j$. In $\mathcal{H}_{2,1}$, the monomial $\alpha_{i}\beta_{j}$ only appears as a coefficient of $X^{(1)}$. Since vectors of the form $X_{i},X_{j}]$ form a basis for $\mathfrak{g}_{1}$, this implies that $\mathcal{H}_{2,1}$ is nondegenerate. From this we conclude the proposition.∎
Next, we show that step-3 Lie algebras are also 2-great. This is the best we can do in terms of step as there exist step-4 Lie algebras that are not 2-great as we shall see. Our proof technique will be similar to the proofs of $2$-greatness that we have already presented in that we begin by fixing some choice of $k$ which we show to be in the complement of $\mathcal{D}(\mathcal{H}_{2,2})$. It will differ in that we will show that the existence of a linear dependence between the polynomials of $\mathcal{H}_{2,2}$ contradicts the fact that $\mathfrak{g}$ is step 3 rather than directly showing linear independence by identifying unique monomials of each polynomial.
@@
Since the equation above holds for arbitrary $i_{1},i_{2},i_{3}\in I$, this implies that $\lambda\!\!=\!\!0$, a contradiction. Thus, $\mathcal{H}_{2,2}$ is non-degenerate. The fact that $\mathcal{H}_{2,0}$ and $\mathcal{H}_{2,1}$ are non-degenerate follows from Proposition 5.3, so any step-3 (or lower) Lie algebra is 2-great. ∎